# PartSDF: Part-Based Implicit Neural Representation for Composite 3D Shape Parametrization and Optimization

### v2

@@ 1 Introduction
To make part-based modeling practical and robust, we develop a decoder architecture that enables dynamic interaction between parts during inference. This decoder involves multiple lightweight convolutions that operate within each part and across parts respectively, enabling part-wise information exchange into a shared intermediate representation and allowing each part to adapt itself based on the geometry of others. This yields consistent deformations, without heavy computations and without resorting to complex hierarchical assemblies or attention-based systems (Deng et al., 2022; O. et al., 2022; Li et al., 2024). Complementing this, we propose a novel supervision scheme that bypasses the need for clean, watertight part meshes. Instead of requiring explicit SDF supervision for each part, we derive supervisory signals from the global shape’s SDF, applied only within regions of space closest to each part. This strategy not only aligns with the natural spatial semantics of parts but also allows effective training on noisy, incomplete, or non-watertight data—significantly broadening the method’s applicability without compromising geometric fidelity.
@@ 2.1 Learned 3D Shape Representation
Learning-based methods for 3D shape representation have evolved significantly, beginning with explicit ones such as voxel grids, point clouds, and meshes. Voxel methods (Wu et al., 2015; 2016; Choy et al., 2016; Dai et al., 2017) partition 3D space into grids but are memory-intensive at finer resolutions, which can be mitigated using octrees (Riegler et al., 2017; Tatarchenko et al., 2017), but only up to a point. Point clouds reduce memory costs by representing shapes as a set of points but ignore connectivity, which can compromise topological consistency (Fan et al., 2017; Yang et al., 2018; Achlioptas et al., 2018; Peng et al., 2021; Zeng et al., 2022). Meshes, though well-suited for detailed surface representation, impose a rigid topology that is difficult to modify (Groueix et al., 2018; Kanazawa et al., 2018; Wang et al., 2018; Pan & Jia, 2019).
@@ 2.2 Part Based Models
Early unsupervised approaches learn shape abstractions using local primitives such as cuboids (Tulsiani et al., 2017; Zou et al., 2017; Sun et al., 2019; Smirnov et al., 2020; Yang & Chen, 2021), superquadrics (Paschalidou et al., 2019; 2020), anisotropic Gaussians (Genova et al., 2019), or convexes (Deng et al., 2020), approximating complex shapes as combinations of simpler parts. More recently, complex objects are better represented by predicting and deforming primitives (Paschalidou et al., 2021; Shuai et al., 2023), creating deformable part templates (Hui et al., 2022), or performing part-based co-segmentation, which consistently divides shapes into parts across a dataset without relying on labeled boundaries (Chen et al., 2023) or does so for only a subset of the data (Chen et al., 2019). Instead, RIM-Net (Niu et al., 2022) learns a hierarchical structure of INRs, while PartNeRF (Tertikas et al., 2023) proposes a rendering-based approach for generating part-aware editable 3D shapes. SPAGHETTI (O. et al., 2022) predicts Gaussian parts from a global latent vector and reconstructs them into a single cohesive object, supporting interactive editing of user-defined parts. It can be combined with SALAD (Koo et al., 2023), which employs a cascaded diffusion model for part generation.
@@ 4 Experiments
We demonstrate our method on several tasks and compare it to multiple part-based baselines. Additional implementation and result details are in Appendices˜B and E.
@@ 4.4 Part Optimization
with respect to $\left(\mathbf{Z},\mathbf{P}\right)$, where $\mathcal{L}_{\text{reg}}$ are optional regularization terms and MC is the Marching Cubes algorithm (Lewiner et al., 2003). Importantly, gradients can be computed through the meshing step (Remelli et al., 2020). In practice, we minimize $\mathcal{L}_{\text{drag}}$ over the PartSDF latents $\mathbf{Z}$ and pose parameters $\mathbf{P}$, starting from their values for an existing car. In the example of Fig. 8, we optimize the shapes of several cars while keeping their wheels unchanged, which is something that our part-based approach allows and can not be done with a global representation such as the one of Ours-1P or 3DShape2VecSet (Zhang et al., 2023). While they may yield a lower final drag, the shapes do not respect design constraints, which is undesirable. For instance, in the second and last row of Figure˜8, results with Ours-1P tend to converge to an average car, ignoring the specificity of the initial cars and their wheels. We provide more details and results in Appendix˜C.
@@ 5 Limitations
The main limitation of PartSDF is its reliance on part labels during training. For our intended application, computer-aided design, this is not a major issue because shapes are typically constructed from individual parts; hence, the decomposition of shapes into parts is known a priori. Thus, PartSDF is particularly relevant in scenarios where objects are naturally created with part-aware structures. While this limits PartSDF’s applicability to datasets without predefined decompositions—such as most of those available online—it better matches with practical engineering requirements. Nonetheless, if generalization to unlabeled data is desired, PartSDF could be combined with co-segmentation methods to generate pseudo-labels for its training. Additionally, while we devised a strategy to handle non-watertight parts, it still assumes that they are volumetric and can be represented by an SDF. Thus, parts that are pure surfaces, such as car hoods, would not be naturally handled without inflating them by wrapping a volume around them. An interesting direction for future work would be to use unsigned distance functions to solve this.
@@ Appendix A Table of Contents
In this supplementary material, we provide details regarding our implementation and experimental setup in Appendix˜B, explain the shape optimization experiment and give additional results in Appendix˜C, show parts interpolation in Appendix˜D, and we provide additional experimental results in Appendix˜E: further visualizations for shape reconstruction (Section˜E.1), generation (Section˜E.2) and manipulation (Section˜E.3).
@@ B.4 Datasets
In order to perform Computational Fluid Dynamics (CFD) simulations on the cars, see Appendix˜C below, we need them to be correctly closed and physically plausible, i.e. wheels and body correctly separated. We therefore hand-process a subset of ShapeNet (Chang et al., 2015)’s cars: The wheels and body are manually selected and separated, and then remeshed to become watertight (Wang et al., 2022).
@@ B.6 Metrics
Appendix C Part Optimization
@@ Results.
with an average relative improvement of 12.7%. We visualize some optimization results in Figure˜10. As can be observed, only the body of the car has been modified while the wheels are unchanged. Nonetheless, they stay consistent with each other and no inter-penetration happens.
Appendix D Part Interpolation
@@
To demonstrate the smoothness and consistency of our latent and pose space, we perform an interpolation experiment. Given pair of shapes, we linearly interpolate between their part latent vectors and poses, with the exception of rotation quaternions for which we use spherical linear interpolation (slerp). At multiple interpolation steps, we reconstruct the intermediate shapes using our decoder and visualize such interpolation "path" in Figure˜11, illustrating smooth transitions between the pair of shapes. Notably, the part-based structure remains coherent throughout the interpolation. This highlights the effectiveness of our learned latent and pose spaces in representing composite shapes.
Appendix E Additional Experimental Results
In this section, we provide additional results and visualization on shape reconstruction (Section˜E.1), generation (Section˜E.2) and manipulation (Section˜E.3).
E.1 Shape Reconstruction
We provide additional examples of shape reconstruction on the test set in Figure˜12. We also visualize the part poses as primitives and results using the point encoder. DAE-Net and BAE-Net struggle to recover thin parts and complex shapes, while PQ-Net tend to inflates them without details. On the opposite, PartSDF recovers accurately the shapes and the parts. We note that with a point cloud encoder, the results may not be as accurate, e.g., the last chair where the armrest segmentation is different, though it is still sensible.
@@ E.1 Shape Reconstruction
E.2 Shape Generation
In Figure˜13, we visualize generation examples with PQ-Net and PartSDF. We also show the part latent generation conditioned on poses for our model. The generated shapes with our method in combination of SALAD (Koo et al., 2023) are more detailed than PQ-Net + a latentGAN, with greater variety for cars and mixers (as reported in Table˜2).

### v3

@@ 1 Introduction
To make part-based modeling practical and robust, we develop a decoder architecture that enables dynamic interaction between parts during inference. This decoder involves multiple lightweight convolutions that operate within each part and across parts respectively, enabling part-wise information exchange into a shared intermediate representation and allowing each part to adapt itself based on the geometry of others. This yields consistent deformations, without heavy computations and without resorting to complex hierarchical assemblies or attention-based systems (Deng et al., 2022; Hertz et al., 2022; Li et al., 2024). Complementing this, we propose a novel supervision scheme that bypasses the need for clean, watertight part meshes. Instead of requiring explicit SDF supervision for each part, we derive supervisory signals from the global shape’s SDF, applied only within regions of space closest to each part. This strategy aligns with the natural spatial semantics of parts and allows effective training even when part segmentations are not watertight, broadening the method’s applicability without compromising geometric fidelity.
@@ 2.1 Learned 3D Shape Representation
Learning-based methods for 3D shape representation have evolved significantly, beginning with explicit ones such as voxel grids, point clouds, and meshes. Voxel methods (Wu et al., 2015; 2016; Choy et al., 2016b; Dai et al., 2017) partition 3D space into grids but are memory-intensive at finer resolutions, which can be mitigated using octrees (Riegler et al., 2017; Tatarchenko et al., 2017), but only up to a point. Point clouds reduce memory costs by representing shapes as a set of points but ignore connectivity, which can compromise topological consistency (Fan et al., 2017; Yang et al., 2018; Achlioptas et al., 2018; Peng et al., 2021; Zeng et al., 2022). Meshes, though well-suited for detailed surface representation, impose a rigid topology that is difficult to modify (Groueix et al., 2018; Kanazawa et al., 2018; Wang et al., 2018; Pan & Jia, 2019).
@@ 2.2 Part Based Models
Early unsupervised approaches learn shape abstractions using local primitives such as cuboids (Tulsiani et al., 2017; Zou et al., 2017; Sun et al., 2019; Smirnov et al., 2020; Yang & Chen, 2021), superquadrics (Paschalidou et al., 2019; 2020), anisotropic Gaussians (Genova et al., 2019), or convexes (Deng et al., 2020), approximating complex shapes as combinations of simpler parts. More recently, complex objects are better represented by predicting and deforming primitives (Paschalidou et al., 2021; Shuai et al., 2023), creating deformable part templates (Hui et al., 2022), or performing part-based co-segmentation, which consistently divides shapes into parts across a dataset without relying on labeled boundaries (Chen et al., 2023) or does so for only a subset of the data (Chen et al., 2019). Instead, RIM-Net (Niu et al., 2022) learns a hierarchical structure of INRs, while PartNeRF (Tertikas et al., 2023) proposes a rendering-based approach for generating part-aware editable 3D shapes. SPAGHETTI (Hertz et al., 2022) predicts Gaussian parts from a global latent vector and reconstructs them into a single cohesive object, supporting interactive editing of user-defined parts. It can be combined with SALAD (Koo et al., 2023), which employs a cascaded diffusion model for part generation.
@@
Part-Aware Generation from Images.
In addition to methods trained with 3D part supervision, several recent works explore part- or object-aware 3D generation, typically from images. Part123 (Liu et al., 2024) and PartGen (Chen et al., 2024) both begin by generating multi-view renderings with associated 2D part segmentations. Part123 then employs NeuS (Wang et al., 2021) to recover part geometries, while PartGen predicts completed part images to fill occlusions before reconstructing their 3D geometry; it can also operate from text prompts in addition to images. Concurrently, MIDI (Huang et al., 2025), addresses single-view scene reconstruction by segmenting objects and conditioning a 3D diffusion model on each object instance, with cross-object attention, while CAST (Yao et al., 2025) performs open-vocabulary scene reconstruction from an RGB image using large-scale pre-trained foundation models, diffusion, and physics-aware post-processing.
Unlike PartSDF, these approaches primarily focus on image-to-3D pipelines and emphasize image reconstruction, as opposed to geometric accuracy or ease of manipulation. In contrast, PartSDF learns supervised, part-aware implicit representations with explicit latent parametrization, enabling coherent part manipulation and constrained optimization, which are crucial in engineering and design scenarios.
@@ 4 Experiments
We demonstrate our method on several tasks and compare it to multiple part-based baselines. Additional implementation details and results are given in Appendices˜B and G.
@@ 4.4 Part Optimization
with respect to $\left(\mathbf{Z},\mathbf{P}\right)$, where $\mathcal{L}_{\text{reg}}$ are optional regularization terms and MC is the Marching Cubes algorithm (Lewiner et al., 2003). Importantly, gradients can be computed through the meshing step (Remelli et al., 2020). In practice, we minimize $\mathcal{L}_{\text{drag}}$ over the PartSDF latents $\mathbf{Z}$ and pose parameters $\mathbf{P}$, starting from their values for an existing car. In the example of Fig. 8, we optimize the shapes of several cars while keeping their wheels unchanged, which is something that our part-based approach allows and can not be done with a global representation such as the one of Ours-1P or 3DShape2VecSet (Zhang et al., 2023). While they may yield a lower final drag, the shapes do not respect design constraints, which is undesirable. For instance, in the second and last row of Figure˜8, results with Ours-1P tend to converge to an average car, ignoring the specificity of the initial cars and their wheels. We provide more details and results in Appendix˜D.
@@ 5 Limitations
The main limitation of PartSDF is its reliance on part labels during training. For our intended application, computer-aided design, this is not a major issue because shapes are typically constructed from individual parts; hence, the decomposition of shapes into parts is known a priori. Thus, PartSDF is particularly relevant in scenarios where objects are naturally created with part-aware structures. While this limits PartSDF’s applicability to datasets without predefined decompositions—such as most of those available online—it better matches with practical engineering requirements. Nonetheless, if generalization to unlabeled data is desired, PartSDF could be combined with co-segmentation methods to generate pseudo-labels for its training. PartSDF also has a few typical failure modes that are illustrated in Fig. 10. These include occasional artifacts when interpolating between shapes with different part existences, residual small surfaces for absent parts, and difficulties representing very thin structures. Such issues occur infrequently and can often be mitigated by discarding known-absent parts or by using higher-resolution meshing for thin geometries. Additionally, because the computational and memory costs scale with the number of parts, extending PartSDF to shapes with high part counts might require strategies to improve efficiency, such as sparse cross-part interactions. Finally, while we devised a strategy to handle non-watertight parts, it still assumes that they are volumetric and can be represented by an SDF. Thus, parts that are pure surfaces, such as car hoods, would not be naturally handled without inflating them by wrapping a volume around them. An interesting direction for future work is to use unsigned distance functions to solve this.
Figure 10: Failures cases of PartSDF. (Left) Interpolation between shapes where a part exists in one but not the other may produce floating artifacts for intermediate shapes (e.g., chair armrests). (Middle) In rare test cases, parts that should be absent can persist as small surface fragments, such as the green armrest patches, which affects Chamfer-Distance. (Right) Very thin structures, such as the helix of some mixers, remain challenging to represent and mesh accurately using an SDF-based approach.
@@ Appendix A Table of Contents
B
Implementation Details
B.1
Network Architecture
B.2
Training Procedure
B.3
Encoder and Diffusion Models
B.4
Datasets
B.5
Baselines
B.6
Metrics
C
Design Analysis
D
Part Optimization
E
Part Interpolation
F
Results on DrivAerNet++
G
Additional Experimental Results
G.1
Shape Reconstruction
G.2
Shape Generation
G.3
Shape Manipulation
H
Single-View Reconstruction
@@ B.4 Datasets
In order to perform Computational Fluid Dynamics (CFD) simulations on the cars, see Appendix˜D below, we need them to be correctly closed and physically plausible, i.e. wheels and body correctly separated. We therefore hand-process a subset of ShapeNet (Chang et al., 2015)’s cars: The wheels and body are manually selected and separated, and then remeshed to become watertight (Wang et al., 2022).
@@ B.5 Baselines
B.5.1 Model size comparison
We compare the models’ size in term of their number of trainable parameters. As can be seen in Table˜4, PartSDF is the smallest model, with 3DShape2VecSet being $\sim 40\times$ larger. In practice, 3DShape2VecSet also requires multi-GPU training, as noted in its paper, while PartSDF can be trained on a single GPU.
Table 4: Model sizes. Size of the different models in term of trainable parameters. *: Most parameters are in the MLPs of the Self-Attention layers between the latent vectors. : We only consider the parameters of the blending network and ignore the object generator, which is only relevant for generation.
# of parameters
3DShape2VecSet* 106.1M
Ours-1P 2.5M
DAE-Net 3.1M
BAE-Net 5.3M
PQ-Net 12.8M
PASTA 9.0M
Ours 2.5M
To understand the good performance of PartSDF in spite of its smaller size, we note that 3DShape2VecSet is an auto-encoding method designed for point cloud reconstruction and diffusion-based generation, not for per-shape optimization. In contrast, PartSDF is an auto-decoding framework that is well-suited for optimization and manipulation tasks. For completeness, we also experiment with optimizing 3DShape2VecSet latents at inference time on the Car dataset, with two settings: "+AD", where latents are randomly initialized and optimized, as in PartSDF, and "+AE+AD", where latents are first obtained from the point-cloud encoder and then further optimized. We report the results in Table˜5.
Table 5: Auto-decoding with 3DShape2VecSet. Reconstruction results when using auto-decoding with 3DShape2VecSet with randomly initialized latents (+AD) or encoded from a point-cloud (+AE+AD).
CD ($\downarrow$) IoU (%, $\uparrow$) IC ($\uparrow$) pIoU (%, $\uparrow$)
3DS2VS 2.73 92.03 0.887 -
3DS2VS+AD 57.97 92.49 0.848 -
3DS2VS+AE+AD 1.15 98.78 0.942 -
Ours-1P 1.37 97.41 0.924 -
Ours 1.28 97.95 0.930 94.25
In the +AD setting, 3DShape2VecSet struggles to optimize its latent vectors and the reconstructions exhibit bumpy surfaces and floaters, hence the lower IC and poor CD. This is consistent with observations made in the DeepSDF paper’s supplementary material (Sec. E, Fig. 19). This highlights that its latent space is not well suited for direct optimization, one of the main reasons we did not adopt a VecSet-style latent space for PartSDF. With +AE+AD, performance improves and slightly surpasses that of PartSDF in reconstruction accuracy.
However, 3DShape2VecSet is not a good fit for the applications we are targeting: It does not provide explicit part representations and its latent space is not designed for controllability or per-part optimization. To demonstrate this, we add small amounts of Gaussian noise to the latent vectors of 3DShape2VecSet and PartSDF before reconstruction. In Figure˜11, the standard deviation of the noise is taken to be 5% and 7.5% of the latents average norm. Adding such small amounts of noise to 3DShape2VecSet’s latent vectors quickly degrades the shape, while PartSDF’s output remains a valid car. In other words, PartSDF’s latent space is better behaved and, thus, better suited for shape optimization.
Figure 11: Reconstruction with noisy latents. When gradually adding gaussian noise to the latent vectors in 3DShape2VecSet and PartSDF, we observe a degradation of the shape with the former. With the latter, the output remains a slightly different, but valid, car. The noise is taken to be 5% (+noise) and 7.5% (++noise) of the latents average norm.
@@ B.6 Metrics
Appendix C Design Analysis
Number of parts.
We evaluate the influence of the number of parts on the Chair dataset by artificially fusing or splitting parts to vary the decomposition. We report the results in Table˜6 and visualize examples of parts decomposition in Figure˜12. Increasing the number of parts generally improves reconstruction metrics, as simpler per-part geometries are easier to represent. However, we observe a slight Chamfer-Distance degradation due to a few failures in removing armrests during reconstruction, which suggests that larger part counts may require more data and training to robustly capture cross-part interactions. Part IoU remains relatively stable, with only minor variation ($\simeq 0.6\%$). Overall, while finer decompositions can improve expressiveness, they may reduce controllability by requiring more poses to be specified and sufficient training data to model all part relationships.
Table 6: Varying the number of parts. Reconstruction results on the Chair dataset with PartSDF when varying the number of parts in the decomposition.
# of parts CD ($\downarrow$) IoU (%, $\uparrow$) IC ($\uparrow$) pIoU (%, $\uparrow$)
1 3.74 93.78 0.913 -
2 1.61 95.91 0.931 93.08
3 2.00 96.55 0.936 93.24
4 1.27 97.13 0.943 93.69
5 1.51 97.23 0.943 93.63
7 1.32 97.21 0.944 93.12
8 1.30 97.17 0.942 93.67
9 1.39 97.37 0.945 93.22
12 1.76 97.48 0.945 93.01
Figure 12: Part decompositions for the Chairs. We show an example of shape with its parts colored differently to visualize the part decompositions.
Architectural variants.
We experiment with a transformer-based architecture (Vaswani et al., 2017) to replace our cross-part decoder with self-attention, see Table˜7. While feasible, this increases parameter count (from $\sim$2.5M to 4.8M), memory (traing samples reduced by a factor 4 to avoid out-of-memory errors) and training cost (from 2000 to 4000 epochs with lower learning-rate) for the same reconstruction accuracy. We also observed greater parts entanglement in this architecture: displacing the car wheels also affect the car length even when its scale is kept constant, see Figure˜13 where the car front and back move with the wheels instead of simply adapting their wells around them. We believe this can be solved, for example by adding some shape mixing during training, but we note that our simpler cross-part decoder already enables it at a lower memory and computation cost.
Table 7: Comparing architectures. Reconstruction results on the Car dataset using a Transformer-based decoder (Transf.) and a softmin SDF-aggregation strategy with our decoder (softmin) against PartSDF.
CD ($\downarrow$) IoU (%, $\uparrow$) IC ($\uparrow$) pIoU (%, $\uparrow$)
Transf. 1.25 98.00 0.931 93.44
softmin 1.31 97.90 0.930 94.78
Ours 1.27 98.02 0.931 94.89
Figure 13: Manipulation with different architectures. We compare manipulation between a Transformer-based variant against our cross-part decoder, by translating the wheels while keeping the body’s scale fixed. For the Transformer-based decoder, the length of the car is incorrectly affected.
SDF aggregation strategies.
The merger of two SDFs into a single one is performed using the $\min$ operator, see, e.g., https://iquilezles.org/articles/distfunctions/, which simply aggregates the parts. We tried replacing the $\min$ by a softmin variant and found that it performs similarly, if marginally worse, as shown in Table˜7 for the Car dataset.
Scalability of PartSDF.
The computational cost of PartSDF depends on both the number of parts $P$ and the feature dimension $D$. The single-part layers $h_{\mathrm{sp}}$ scale as $O(PD^{2})$, while the cross-part layers $h_{\mathrm{cp}}$ scale as $O(P^{2}D)$. In our experiments, $D\in\{256,512\}$ and $P\ll D$, so runtime and memory scale approximately linearly with the number of parts. However, for very large $P$, the combined complexity becomes $O(PD^{2}+P^{2}D)$, similar to the cost of transformer layers (considering self-attention + MLP). Although our layers remain more lightweight, modeling highly complex objects with many parts would require both large $D$ and $P$ to capture fine geometric details, which may limit scalability. Extending our work to such cases is possible but will require strategies such as positional encoding, hierarchical part grouping, or sparse cross-part interactions to improve efficiency.
Appendix D Part Optimization
@@ Results.
with an average relative improvement of 12.7%. We visualize some optimization results in Figure˜14. As can be observed, only the body of the car has been modified while the wheels are unchanged. Nonetheless, they stay consistent with each other and no inter-penetration happens.
Appendix E Part Interpolation
@@
To demonstrate the smoothness and consistency of our latent and pose space, we perform an interpolation experiment. Given pair of shapes, we linearly interpolate between their part latent vectors and poses, with the exception of rotation quaternions for which we use spherical linear interpolation (slerp). At multiple interpolation steps, we reconstruct the intermediate shapes using our decoder and visualize such interpolation "path" in Figure˜15, illustrating smooth transitions between the pair of shapes. Notably, the part-based structure remains coherent throughout the interpolation. This highlights the effectiveness of our learned latent and pose spaces in representing composite shapes.
Appendix F Results on DrivAerNet++
To evaluate PartSDF in a realistic industrial setting, we trained our model on DrivAerNet++ (Elrefaie et al., 2024), a high-fidelity automotive dataset designed for aerodynamic analysis and shape optimization. The dataset includes three base car designs, each with thousands of geometric variations capturing diverse body, back, and wheel configurations.
We used these data to train PartSDF with separate parts for the car body, car back, and wheels, using the same architecture and loss functions as in our main experiments. Figure 16 shows reconstruction examples on held-out test shapes, demonstrating that PartSDF accurately captures fine geometric details and maintains part consistency on these industrial-grade models.
Furthermore, as illustrated in Figure 17, PartSDF enables smooth and coherent manipulation and interpolation between car designs. These results highlight the ability of PartSDF to generalize to high-fidelity shapes and support structured manipulation in practical aerodynamic design scenarios.
Figure 16: Reconstructions on DrivAerNet++. Reconstructions from the test set of the DrivAerNet++ dataset. PartSDF successfully captures fine geometric details and preserves consistency between parts.
Figure 17: Manipulation on DrivAerNet++. (Left) Wheels are translated, then the car body is elongated. (Right-top) We interpolate from an Estateback to a Notchback in a smooth manner. (Right-bottom) We sample random body and back latents with the same wheels and poses to generate varying designs with similar general scale.
Appendix G Additional Experimental Results
In this section, we provide additional results and visualization on shape reconstruction (Section˜G.1), with new object categories, generation (Section˜G.2) and manipulation (Section˜G.3).
G.1 Shape Reconstruction
We provide additional examples of shape reconstruction on the test set in Figure˜18. We also visualize the part poses as primitives and results using the point encoder. DAE-Net and BAE-Net struggle to recover thin parts and complex shapes, while PQ-Net tend to inflates them without details. On the opposite, PartSDF recovers accurately the shapes and the parts. We note that with a point cloud encoder, the results may not be as accurate, e.g., the last chair where the armrest segmentation is different, though it is still sensible.
@@ G.1 Shape Reconstruction
Additionally, we expand our evaluation to additional categories from ShapeNet and PartNet: Display and Knife, featuring 640 and 324 shapes with 2 and 3 parts respectively. Results are given in Table˜8 and Figure˜19. They confirm that PartSDF delivers a strong performance beyond the original categories. In particular, the CD value for our method using multiple parts is markedly better on the Displays, mostly because baselines tend to fail drastically for a few shapes, whereas our complete approach succeeds. Voxel-based approaches DAE-Net, BAE-Net, and PQ-Net particularly struggle on thinner shapes, which can result in very large CDs. We also observe that PASTA struggles on all datasets because it only uses parts bounding box as information for reconstruction.
Table 8: Additional categories. We report reconstruction metrics on test shapes for two new categories: Display and Knife.
CD ($\downarrow$) IoU (%, $\uparrow$) IC ($\uparrow$) pIoU (%, $\uparrow$)
Display Knife Display Knife Display Knife Display Knife
3DShape2VecSet 4.95 1.72 92.57 80.33 0.949 0.870 - -
Ours-1P 4.45 0.35 95.02 93.89 0.949 0.943 - -
DAE-Net 98.75 489.58 58.67 46.63 0.772 0.567 45.76 42.39
BAE-Net 161.16 387.71 43.72 34.26 0.718 0.536 29.81 31.41
PQ-Net 95.73 41.60 42.81 30.45 0.775 0.560 41.44 36.09
PASTA 148.07 10.57 60.01 58.01 0.834 0.718 - -
Ours 1.73 0.28 97.37 96.16 0.965 0.958 89.11 92.81
Figure 19: Additional categories. Reconstructions of test shapes using the baselines and our model on two new categories: Display and Knife.
G.2 Shape Generation
In Figure˜20, we visualize generation examples with PQ-Net and PartSDF. We also show the part latent generation conditioned on poses for our model. The generated shapes with our method in combination of SALAD (Koo et al., 2023) are more detailed than PQ-Net + a latentGAN, with greater variety for cars and mixers (as reported in Table˜2).